For each of the Z-scores below, determine two probabilities (show four three decimal places): (1) the probability above the Z-score and (2) the probability below the Z-score

A. 2.40:

Probability above:

Probability below:

B. -0.48:

Probability above:
  
Probability below:
Answer #1

solution:

p(z>2.40)=

1 - P(z <2.40 )

Using z table

= 1 - 0.9918

probability= 0.0082

below=

P(z <2.40 ) =0.9918

b.

p(z>-0.48)=

1 - P(z <-0.48 )

Using z table

= 1 - 0.3156

probability= 0.6844

below=

P(z <-0.48 ) =0.3156